OpenStudy (anonymous):

A company manufacturers and sells x electric drills per month. The monthly cost and price-demand equations are C(x)=64000+60x, p=190−x/30, 0≤x≤5000. a) production level at max revenue = 2850 b) price to max profit = $125 c) Suppose that a 5 dollar per drill tax is imposed. Determine the number of drills that should be produced and sold in order to maximize profit under these new circumstances. ???

OpenStudy (anonymous):

how did you get the answers?

OpenStudy (anonymous):

R(x) = P(x) * x = (190−x/30)*x

OpenStudy (anonymous):

to find the max revenue, take the derivative and set it equal to zero; solve for x.

OpenStudy (anonymous):

that's the number of units that give max revenue. call it x1. P(x1) is the price at max revenue...

OpenStudy (anonymous):

profit is R(x) -C(x) ...

OpenStudy (anonymous):

same thing, take the derivative and find the max. ... find the price at that production level...